# Characterization of Deletion/Substitution Channel Capacity for Small Deletion and Substitution Probabilities

### v2

@@ Abstract
We consider binary input deletion/substitution channels, which model certain channels with synchronization errors encountered in practice. Specifically, we focus on the regime of small deletion and substitution probabilities, and by extending an approach developed for the deletion-only channel, we obtain an asymptotic characterization of the channel capacity for independent and identically distributed deletion/substitution channels. We first present an upper bound on the capacity for arbitrary but fixed numbers of deletions and substitutions, and then we extend the result to the case of random deletions and substitutions. Our final result is as follows: The i.i.d. deletion/substitution channel capacity is approximately $1-H(p_{d})-H(p_{s})$, for $p_{d},p_{s}\approx 0$, where $p_{d}$ is the deletion probability, and $p_{s}$ is the substitution probability.
@@ I Introduction
Channels with insertions, deletions, and substitutions are crucial in modeling certain communication systems with synchronization errors. For instance, when there is a clock mismatch in a wireless communication system, such errors are observed. In addition, bit-patterned media recording channels and DNA data storage channels exhibit such errors. In this paper, our focus is on the special case of channels experiencing bit losses (deletions) and bit errors (substitutions). Specifically, we are interested in independent, identically distributed (i.i.d.) deletions and substitutions at the regime of small deletion and substitution probabilities. Our objective is to characterize the channel capacity for this asymptotic regime.
Dobrushin established the information stability of channels with memoryless synchronization errors in [1], which is extended in [2] to synchronization errors governed by a Markov chain. However, obtaining the exact capacity has proved very difficult. Even for a channel with only i.i.d. deletions, the capacity is still unknown. As a result, the literature has focused on obtaining bounds on the capacity. Gallager’s work on sequential decoding for channels with synchronization errors provides the first lower bound on the capacity of a deletion channel [3], which states that for deletion probabilities of less than $1/2$, i.e., $p_{d}<1/2$, the capacity of i.i.d. binary deletion channel (BDC) is not smaller than that of a binary symmetric channel with a cross-over probability of $p_{d}$; that is,
@@
where $H(\cdot)$ denotes the binary entropy function. Indeed, Gallager’s work is more general and establishes a lower bound on the capacity of i.i.d. insertion, deletion, and substitution error channels. For the case of deletion/substitution channels (i.e., with insertion probability of zero), the lower bound becomes:
@@
where $p_{s}$ is the substitution probability, and $p_{c}=1-p_{d}-p_{s}$ is the probability of correct transmission. In the regime of asymptotically small synchronization errors, this can be lower bounded as
$C_{d,s}\geq 1-H(p_{d})-H(p_{s}).$ (3)
We will prove that this bound is asymptotically tight by establishing an upper bound on the capacity whose dominant terms match it in the asymptotic regime of $p_{d},p_{s}\rightarrow 0$. There are also other attempts to obtain lower bounds on the capacity of channels with synchronization errors. The authors in [4] employ tree codes to obtain lower bounds for insertion/deletion channels. In [5], a lower bound on the BDC capacity is obtained using the jigsaw-puzzle decoding approach, which is further improved in [6].
Only a limited number of upper bounds have been presented for the capacity of channels with synchronization errors. In [7], the authors provide two upper bounds on the BDC capacity: a numerically evaluated one for any arbitrary deletion probability and a simple closed-form one. In [8], numerical upper bounds are provided for deletion/substitution channels by exploiting an auxiliary system where the receiver is provided with side information. The authors extend their results to a non-binary deletion channel in [9], and to more general synchronization error channels with insertions, deletions, and substitutions in [10]. The authors of [11] improve the BDC lower bound in [6] as well as the upper bound in [8]. In [12], by converting the capacity problem into maximizing a univariate, real-valued, and often concave function, the authors obtain upper bounds for a general class of repetition channels.
Several works have focused on approximating the capacity of synchronization error channels for the asymptotic regime of small synchronization errors. In [13, 14], the BDC capacity is approximated using series expansions up to the second and third leading terms, respectively, by calculating the dominant terms of the capacity expansion for small deletion probabilities. Similar results are obtained for duplication channels [15] and insertion channels [16].
More in line with our work, the authors of [17] show that, for small deletion probabilities, the capacity can be upper-bounded as
@@
To obtain this bound, they first obtain lower and upper bounds on the success probability of a guesser that, given the received erroneous sequence, outputs the transmitted sequence along with the deletion pattern. To this end, assuming that the decoding is successful with a certain probability, they obtain bounds on the number of deletion patterns in a certain vicinity of the desired one that can be mistaken for it at the receiver. Finally, combining (4) with Gallager’s lower bound, they show that the BDC capacity approaches $1-H(p_{d})$ as $p_{d}\rightarrow 0$. The authors extend the result to the $k$-copy BDC in [18].
In this paper, we adapt the general methodology in [17] and extend it to channels with substitution errors as well as deletions. We first provide an upper bound on the capacity of deletion/substitution channels. To this end, we first fix the numbers of deletions and substitutions and assume that the receiver knows these numbers as side information (it knows the number of deletions anyway due to the received sequence length), and obtain an upper bound on the capacity, which is also an upper bound on the capacity of the case without side information since side information cannot decrease capacity. Combining with the result in [3], we obtain a tight approximation of the capacity when the deletion and substitution probabilities are asymptotically small. Finally, we extend the result to the case of random numbers of deletions and substitutions, and obtain a tight approximation for the capacity in the asymptotic regime of small deletion and substitution probabilities.
@@ III Fixing the numbers of deletions and substitutions
We aim to obtain an upper bound on the size of a uniquely decodable codebook, whose logarithm gives us an upper bound on the capacity. We first obtain the upper bound on the codebook size for fixed numbers of deletions and substitutions, assuming that the receiver knows them as side information. Note that the receiver is always aware of the number of deletions, as it is equal to the length loss of the received sequence, but not the number of substitutions.
@@
Lemma 1 ([17])
Let $\rho$ be a joint distribution over $S\times T$, for finite sets $S$ and $T$, such that the marginal distribution over $S$ is uniform. Let $g:T\rightarrow S$ be a function. Then, for $(a,b)$ following the joint distribution $\rho$, $Pr[g(b)=a]\leq\frac{|T|}{|S|}$.
@@
To obtain a lower bound on the success probability, we first need to obtain an upper bound on the number of neighboring patterns that can be mistaken for the actual one. To this end, we first upper bound the probability of two transmission/substitution pattern pairs leading to the same output for all possible inputs in the following lemma.
@@
Take any two transmission/substitution pattern pairs, $A=(A_{t},A_{s})$ and $B=(B_{t},B_{s})$. For uniformly random $X\in\{0,1\}^{n}$, $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$ if $\delta_{s}\subseteq\Delta_{t}$, and zero otherwise, where $\Delta_{t}$ denotes the discrepancy set between the two transmission patterns and $\delta_{s}$ is the symmetric difference of the two substitution patterns [^1] Note that $\Delta_{t}$ is defined over the indices of the elements of the two sequences, not on their values.; i.e., $\Delta_{t}=\Delta(A_{t},B_{t})$ and $\delta_{s}=\delta(A_{s},B_{s})$.
@@
In [17], it is shown that in the case of only deletions, the probability of channel outputs being the same is upper bounded as $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A_{t}}=X_{B_{t}}]\leq 2^{-|\Delta_{t}|}$. In the case of only substitutions, two outputs can be the same only if the two substitution patterns are exactly the same. However, that is not the case for the deletion/substitution scenario. In that case, as long as the substitution pattern difference set is a subset of the discrepancy set of the transmission patterns, the two outputs can be the same with the same probability as the two transmission patterns that lead to the same output. So, we have
$\begin{split}Pr_{X\in U^{\{0,1\}^{n}}}&[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\\ &\begin{cases}=1,&\text{if }A=B\\ =0,&\text{if }A_{t}=B_{t},A_{s}\neq B_{s}\\ \leq 2^{-|\Delta_{t}|},&\text{if }A_{t}\neq B_{t},A_{s}=B_{s}\\ =0,&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\not\subseteq\Delta_{t}\\ \leq 2^{-|\Delta_{t}|}.&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\subseteq\Delta_{t}\\ \end{cases}\end{split}$ (6)
@@
The expression in Lemma 1 can be further upper bounded as $Pr_{X\in U^{\{}0,1\}^{n}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$.
In the next lemma, we upper bound the number of patterns in a certain neighborhood of a given pattern that can be mistaken for it.
@@
$\begin{split}&I=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|\leq t,|\delta_{s}|\leq s\}|\\ &\leq\sum_{t^{\prime}=0}^{t}\underbrace{|\{B||\Delta_{t}|=t^{\prime}\}|}_{I_{t^{\prime}}}\times\underbrace{|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|}_{I_{s}}.\end{split}$ (8)
$I_{t^{\prime}}$ is upper bounded in [17] as
$I_{t^{\prime}}\leq{{2q_{d}+t^{\prime}+1}\choose{2q_{d}+1}}{{q_{d}+t^{\prime}}\choose{q_{d}}}$ (9)
@@
$\begin{split}I_{s}&=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|\\ &=\sum_{s^{\prime}=0}^{s}|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|=s^{\prime}\}|\\ &\leq\sum_{s^{\prime}=0}^{s}{q_{s}\choose q_{s}-s^{\prime}}{t^{\prime}\choose s^{\prime}},\end{split}$ (10)
@@
Definition: For integers $t$ and $s$, we say that $X\in\{0,1\}^{n}$ is $(t,s)$-bad if there exist two patterns $A$ and $B$ such that $|\Delta_{t}|\geq t$ or $|\delta_{s}|\geq s$, and $X_{A}=X_{B}$.
@@
We first obtain the probability that a sequence is $(t,s)$-bad, $P_{(t,s)-bad}$. According to Corollary 2, the probability that two patterns $A$ and $B$ lead to the same output is upper bounded by $2^{-|\Delta_{t}|}$. Since this probability is decreasing in $|\Delta_{t}|$, and using union bound trick, we have
$P_{(t,s)-bad}$ $=\!Pr_{X\in U^{\{0,1\}^{n}}}[\exists A,B|(|\Delta_{t}|\!\geq t\lor\!|\delta_{s}|\geq s)\!\land\!X_{A}\!\!=\!\!X_{B}]$
$\leq\sum_{A,B}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}]$
($\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{-t}$ (14)
Finally, multiplying this probability by the total number of sequences concludes the proof. ∎
@@
Fixing the numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta\geq 0$. Then, the size of the codebook satisfies
@@
where
$\alpha=t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1},$
with $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil.$
@@
We first assume that the receiver knows the numbers of deletions and substitutions as side information. We consider a guesser that, given the received sequence $Z_{A}$, outputs the transmitted sequence $Z$ and the pattern pair $A=(A_{t},A_{s})$, both chosen uniformly at random, with a non-negligible probability. More specifically, on the input $Z_{A}$, the guesser first decodes it with the proposed decoder $R(\cdot)$ with success probability $\delta$, then outputs $g(Z_{A})=(R(Z_{A}),A^{\prime})$, where $A^{\prime}$ is the lexicographically first sequence to satisfy $(R(Z_{A}))_{A^{\prime}}=Z_{A}$ if one exists, and the fixed pattern pair $A^{\prime}=(A_{t}^{\prime}=1,2,\cdots,n-q_{d},A_{s}^{\prime}=1,2,...,q_{s})$ otherwise.
We now lower bound the success probability of the guesser. For this, we first consider the probability that the decoding succeeds ($R(Z_{A})=Z$) and the codeword $Z\in\mathcal{C}$ is not $(t,s)$-bad as follows
@@
If $N\leq\frac{2^{n}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, the theorem follows trivially. For $N\geq\frac{2^{n}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, we have
$\begin{split}\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}&\leq 2^{-t}{n\choose q_{d}}^{3}{n-q_{d}\choose q_{s}}^{3}\\ &\leq 2^{-t}\left(\frac{ne}{q_{d}}\right)^{3q_{d}}\left(\frac{(n-q_{d})e}{q_{s}}\right)^{3q_{s}},\end{split}$ (17)
making which smaller or equal to $\frac{\delta}{2}$ (our own choice), gives us a lower bound on $t$ as $t\geq{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$. We take $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil$.
If $Z$ is not $(t,s)$-bad, pattern $B$ gets mixed up with the desired pattern $A$ only when $\Delta_{t}\leq t-1$ and $\delta_{s}\leq s-1$. According to Lemma 7, the number of such patterns is upper bounded by
@@
knowing that $s\leq t$ due to $\delta_{s}\subseteq\Delta_{t}$ for two patterns to be mistaken with each other, and using the fact that (${n\choose k}\leq\left(ne/k\right)^{k}$. We also used $t\geq 3q_{d}$ and $t\geq 3q_{s}$, since $t\geq{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$.
@@
Since side information cannot decrease the capacity, this is also an upper bound on the codebook size of the case without side information. This concludes the proof. ∎
IV Random numbers of deletions and substitutions
@@
In the setup with random numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta$ for codeword length of $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$. Then, the codebook size $N$ satisfies
$\begin{split}\log N\leq n-q_{d,\min}-\log{n\choose q_{d,\min}}&-\log{n-q_{d,\max}\choose q_{s,\min}}\\ &+\log\frac{8}{\delta^{2}}+\log\alpha_{\max}^{*},\end{split}$ (22)
where $q_{d,\min}=(1-\gamma)np_{d}$, $q_{s,\min}=(1-\gamma)n(1-p_{d})p_{s}^{\prime}$, $q_{d,\max}=(1+\gamma)np_{d}$, $q_{s,\max}=(1+\gamma)n(1-p_{d})p_{s}^{\prime}$, with Let $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$, and
$\alpha_{\max}^{*}={t_{\max}^{*2}}\left(2e\right)^{t_{\max}^{*}-1}\left(\frac{5t_{\max}^{*}}{q_{s,\max}^{*}}\right)^{q_{s,\max}^{*}}\left(\frac{5t_{\max}^{*}}{q_{d,\max}^{*}}\right)^{3q_{d,\max}^{*}+1},$ (23)
with $t_{\max}^{*}=\lceil{3q_{d,\max}^{*}}\log\frac{ne}{q_{d,\max}^{*}}+{3q_{s,\max}^{*}}\log\frac{e(n-q_{d,\max}^{*})}{q_{s,\max}^{*}}+\log\frac{8}{\delta^{2}}\rceil$.

### v3

@@ Abstract
We consider binary input deletion/substitution channels, which model certain types of synchronization errors encountered in practice. Specifically, we focus on the regime of small deletion and substitution probabilities, and by extending an approach developed for the deletion-only channel, we obtain an asymptotic characterization of the channel capacity for independent and identically distributed (i.i.d.) deletion/substitution channels. To do so, given a target probability of successful decoding, we first develop an upper bound on the codebook size for arbitrary but fixed numbers of deletions and substitutions, and then extend the result to the case of random deletions and substitutions to obtain a bound on the channel capacity. Our final result is: The i.i.d. deletion/substitution channel capacity is approximately $1-H(p_{d})-H(p_{s})$, for $p_{d},p_{s}\approx 0$, where $p_{d}$ and $p_{s}$ are the deletion and substitution probabilities, respectively.
@@ I Introduction
Channels with insertions, deletions, and substitutions are crucial in modeling certain communication systems with synchronization errors. Such errors are observed in cases like clock mismatches in wireless communication systems, as well as in bit-patterned media recording and DNA data storage channels. In this paper, our focus is on the special case of channels experiencing bit losses (deletions) and bit errors (substitutions). Specifically, we are interested in independent and identically distributed (i.i.d.) deletions and substitutions at the regime of small deletion and substitution probabilities. Our objective is to characterize the channel capacity for this asymptotic regime.
Dobrushin established the information stability of channels with memoryless synchronization errors in [1], which is extended in [2] to synchronization errors governed by a Markov chain. However, obtaining the exact capacity has proved very difficult. Even for a channel with only i.i.d. deletions, the capacity remains unknown. As a result, the literature has focused on obtaining bounds on the capacity. Gallager’s work on sequential decoding for channels with synchronization errors provides the first lower bound on the capacity of a deletion channel [3], which states that for deletion probabilities of less than 1/21/21 / 2, i.e., $p_{d}<1/2$, the capacity of i.i.d. binary deletion channel (BDC) is not smaller than that of a binary symmetric channel with a cross-over probability of $p_{d}$; that is, the i.i.d. deletion channel capacity, denoted by $C_{d}$, satisfies
@@
where $H(\cdot)$ denotes the binary entropy function. Indeed, Gallager’s work is more general and establishes a lower bound on the capacity of i.i.d. insertion, deletion, and substitution error channels. For the case of deletion/substitution channels (i.e., with insertion probability of zero), the lower bound on the capacity, denoted by $C_{d,s}$, becomes:
@@
where $p_{s}$ is the substitution probability, and $p_{c}=1-p_{d}-p_{s}$ is the probability of correct transmission. In the regime of asymptotically small synchronization errors, the lower bound is approximately $1-H(p_{d})-H(p_{s})$.
We will prove that this bound is asymptotically tight by establishing an upper bound on the capacity whose dominant terms match with it in the asymptotic regime of $p_{d},p_{s}\rightarrow 0$. There are also other attempts to obtain lower bounds on the capacity of channels with synchronization errors. The authors in [4] employ tree codes to obtain lower bounds for insertion/deletion channels. In [5], a lower bound on the BDC capacity is obtained using the jigsaw-puzzle decoding approach, which is further improved in [6].
In [7], the authors provide two upper bounds on the BDC capacity: a numerically evaluated one for any arbitrary deletion probability and a simple closed-form one. In [8], numerical upper bounds are provided for deletion/substitution channels by exploiting an auxiliary system where the receiver is provided with side information. The authors extend their results to a non-binary deletion channel in [9], and to more general synchronization error channels with insertions, deletions, and substitutions in [10]. The authors of [11] improve the BDC lower bound in [6] as well as the upper bound in [8]. In [12], by converting the capacity problem into maximization of a real-valued, often concave, univariate function, the authors obtain upper bounds for a general class of repetition channels.
Several works have focused on approximating the capacity of synchronization error channels for the asymptotic regime of small synchronization error probabilities. In [13, 14], the BDC capacity is approximated using series expansions up to the second and third leading terms, respectively. Similar results are obtained for duplication channels in [15] and insertion channels in [16].
More in line with our work, the authors of [17] show that, for small deletion probabilities, the i.i.d. BDC capacity can be upper-bounded as
@@
To obtain this bound, they first obtain lower and upper bounds on the success probability of a guesser that, given the received erroneous sequence, outputs the transmitted sequence along with the deletion pattern. To this end, assuming that the decoding is successful with a certain probability, they obtain bounds on the number of deletion patterns that can be mistaken for the actual one at the receiver. Finally, combining (3) with Gallager’s lower bound, they show that the BDC capacity approaches $1-H(p_{d})$ as $p_{d}\rightarrow 0$. The authors extend the result to the $k$-copy BDC in [18].
In this paper, we adapt the general methodology in [17], and extend the results to channels with substitution errors as well as deletions. Given a probability of successful decoding, we first provide an upper bound on the codebook size for deletion/substitution channels. To this end, we first fix the numbers of deletions and substitutions and assume that the receiver knows these numbers as side information (it knows the number of deletions anyway due to the received sequence length), and obtain an upper bound on the codebook size, which is also an upper bound on the codebook size of the case without side information since side information cannot decrease the number of possible codewords. Finally, we extend the result to the case of random numbers of deletions and substitutions, and, combining with the result in [3], we obtain a tight approximation for the channel capacity in the asymptotic regime of small deletion and substitution probabilities by letting the codeword length go to infinity.
@@ III Fixing the Numbers of Deletions and Substitutions
We aim to obtain an upper bound on the size of a uniquely decodable codebook, given a probability of successful decoding. We first obtain the upper bound on the codebook size for fixed numbers of deletions and substitutions, assuming the receiver knows these values as side information. Note that the receiver is always aware of the number of deletions, as it is equal to the length loss of the received sequence, but not the number of substitutions.
@@
Lemma 1 (Taken from [17])
Let $\rho$ be a joint distribution over $S\times T$, for finite sets $S$ and $T$, such that the marginal distribution over $S$ is uniform. Let $g:T\rightarrow S$ be a function. Then, for $(a,b)\sim\rho$, $Pr[g(b)=a]\leq\frac{|T|}{|S|}$.
@@
To obtain a lower bound on the success probability, we first need to obtain an upper bound on the number of neighboring patterns that can be mistaken for the actual one. To this end, we first upper bound the probability of two transmission/substitution pattern pairs leading to the same output for all possible inputs in the following lemma by extending the result in [17] to the case of both deletion and substitution errors.
@@
Take any two transmission/substitution pattern pairs, $A=(A_{t},A_{s})$ and $B=(B_{t},B_{s})$. For uniformly random $X\in\{0,1\}^{n}$, $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}|A,B]\leq 2^{-|\Delta_{t}|}$ if $\delta_{s}\subseteq\Delta_{t}$, and zero otherwise, where $\Delta_{t}$ denotes the discrepancy set between the two transmission patterns and $\delta_{s}$ is the symmetric difference of the two substitution patterns; i.e., $\Delta_{t}=\Delta(A_{t},B_{t})$ and $\delta_{s}=\delta(A_{s},B_{s})$. Note that $\Delta_{t}$ is defined over the indices of the elements of the two sequences, not on their values.
@@
In [17], it is shown that in the case of only deletions, the probability of channel outputs being the same is upper bounded as $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A_{t}}=X_{B_{t}}]\leq 2^{-|\Delta_{t}|}$. In the case of only substitutions, two outputs can be the same only if the two substitution patterns are exactly the same. However, that is not the case for the deletion/substitution scenario. This is because, as long as the substitution pattern difference set is a subset of the discrepancy set of the transmission patterns, the two outputs can be the same with the same probability as the two transmission patterns that lead to the same output; i.e., $Pr[X_{A}=X_{B}|\delta_{s}\subseteq\Delta_{t}]=Pr[X_{A_{t}}=X_{B_{t}}]$. We obtain
$\begin{split}Pr_{X\in U^{\{0,1\}^{n}}}&[X_{A}=X_{B}|A,B]\\ &\begin{cases}=1,&\text{if }A=B,\\ =0,&\text{if }A_{t}=B_{t},A_{s}\neq B_{s},\\ \leq 2^{-|\Delta_{t}|},&\text{if }A_{t}\neq B_{t},A_{s}=B_{s},\\ =0,&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\not\subseteq\Delta_{t},\\ \leq 2^{-|\Delta_{t}|},&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\subseteq\Delta_{t}.\\ \end{cases}\end{split}$ (5)
@@
The expression in Lemma 2 can be further upper bounded as $Pr_{X\in U^{\{}0,1\}^{n}}[X_{A}=X_{B}|A,B]\leq 2^{-|\Delta_{t}|},\forall\delta_{s}$.
In the following lemma, we upper bound the number of patterns in a certain neighborhood of a given pattern that can be mistaken for it.
@@
$\begin{split}&I=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|\leq t,|\delta_{s}|\leq s\}|\\ &=\!\sum_{t^{\prime}=0}^{t}\underbrace{|\{B_{t}||\Delta_{t}|=t^{\prime}\}|}_{I_{t^{\prime}}}\!\times\!\underbrace{|\{B_{s}|\delta_{s}\!\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\!\leq s\}|}_{I_{s}}.\end{split}$ (7)
$I_{t^{\prime}}$ is upper bounded in [17] as $I_{t^{\prime}}\leq{{2q_{d}+t^{\prime}+1}\choose{2q_{d}+1}}{{q_{d}+t^{\prime}}\choose{q_{d}}}$.
@@
$\begin{split}I_{s}&=|\{B_{s}|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|\\ &=\sum_{s^{\prime}=0}^{s}|\{B_{s}|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|=s^{\prime}\}|\\ &\leq\sum_{s^{\prime}=0}^{s}{q_{s}\choose q_{s}-s^{\prime}}{t^{\prime}\choose s^{\prime}},\end{split}$ (8)
@@
[…]
@@
We first obtain the probability that a randomly picked sequence is $(t,s)$-bad, denoted by $P_{(t,s)-bad}$. According to Corollary 2, the probability that two patterns $A$ and $B$ lead to the same output is upper bounded by $2^{-|\Delta_{t}|}$, which is decreasing in $|\Delta_{t}|$. Then, using the union bound trick, we have
$\begin{split}P_{(t,s)-bad}&\leq\sum_{A,Bs.t.|\Delta_{t}|\geq t}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}]\\ &\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{-t}.\end{split}$ (12)
[…]
@@
Fixing the numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta>0$. Then, the size of the codebook satisfies
@@
where $\alpha=t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1}$, with $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil.$
@@
We first assume that the receiver knows the numbers of deletions and substitutions as side information. We consider a guesser that, given the received sequence $Z_{A}$, outputs the transmitted sequence $Z$ and the pattern pair $A=(A_{t},A_{s})$, both chosen uniformly at random, with a non-negligible probability. More specifically, the guesser first decodes the input $Z_{A}$ with the proposed decoder $R(\cdot)$ with success probability $\delta$, then outputs $g(Z_{A})=(R(Z_{A}),A^{\prime})$, where $A^{\prime}$ is the lexicographically first sequence to satisfy $(R(Z_{A}))_{A^{\prime}}=Z_{A}$ if one exists, and the fixed pattern $A^{\prime}=(A_{t}^{\prime},A_{s}^{\prime})$, with $A_{t}^{\prime}=1,2,\cdots,n-q_{d}$ and $A_{s}^{\prime}=1,2,...,q_{s}$, otherwise.
[…]
@@
If $N\leq\frac{2^{n-q_{d}}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, the theorem follows trivially. For $N\geq\frac{2^{n-q_{d}}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, we have
$\begin{split}\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}&\leq 2^{q_{d}-t}{n\choose q_{d}}^{3}{n-q_{d}\choose q_{s}}^{3}\\ &\leq 2^{q_{d}-t}\left(\frac{ne}{q_{d}}\right)^{3q_{d}}\left(\frac{(n-q_{d})e}{q_{s}}\right)^{3q_{s}}.\end{split}$ (15)
The right hand side is made smaller than or equal to $\frac{\delta}{2}$ if $t$ is selected as $t\geq q_{d}+{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$. We take $t=\lceil q_{d}+{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil$, which lower bounds the right hand side of (15) by $\frac{\delta}{2}$.
[…]
@@
which is obtained by knowing that $s\leq t$ due to $\delta_{s}\subseteq\Delta_{t}$ for two patterns to be mistaken with each other, and using the fact that (${n\choose k}\leq\left(ne/k\right)^{k}$. We also used $t\geq 3q_{d}$ and $t\geq 3q_{s}$, since $t\geq q_{d}+{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$.
@@
Since side information cannot decrease the codebook size, this also serves as an upper bound on the codebook size for the case without side information. This concludes the proof. ∎
IV Capacity of I.I.D. Deletion/Substitution Channel
@@
In the setup with random numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta$ for codeword length of $n\geq\frac{12\log\frac{8}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$. Then, the codebook size $N$ satisfies
$\begin{split}\log N\leq n-q_{d,\min}-\log{n\choose q_{d,\min}}&-\log{n-q_{d,\max}\choose q_{s,\min}}\\ &+\log\frac{4}{\delta}+\log\alpha_{\max}^{*},\end{split}$ (20)
where $q_{d,\min}\!=\!(1\!-\!\gamma)np_{d}$, $q_{s,\min}\!=\!(1\!-\!\gamma)n(1\!-\!p_{d})p_{s}^{\prime}$, $q_{d,\max}\!=\!(1\!+\!\gamma)np_{d}$, $q_{s,\max}\!=\!(1\!+\!\gamma)n(1\!-\!p_{d})p_{s}^{\prime}$, with $\gamma\!=\!\sqrt{\frac{3\log\frac{8}{\delta}}{np_{d}}}$, and $\alpha_{\max}^{*}\!=\!{t_{\max}^{*2}}\left(2e\right)^{t_{\max}^{*}-1}\left(\frac{5t_{\max}^{*}}{q_{s,\max}^{*}}\right)^{q_{s,\max}^{*}}\left(\frac{5t_{\max}^{*}}{q_{d,\max}^{*}}\right)^{3q_{d,\max}^{*}+1}$, with $t_{\max}^{*}=\lceil q_{d,\max}^{*}+{3q_{d,\max}^{*}}\log\frac{ne}{q_{d,\max}^{*}}+{3q_{s,\max}^{*}}\log\frac{e(n-q_{d,\max}^{*})}{q_{s,\max}^{*}}+\log\frac{4}{\delta}\rceil$.